OpenStudy (anonymous):

Translate the following phrases into algebraic expressions. Explain, in complete sentences, the difference between the two. (3 points) •twice the sum of x and y •the sum of twice x and y

hero (hero):

Twice the sum of x and y = 2(x+y) The sum of twice x and y = 2x + y

hero (hero):

Get it?

OpenStudy (anonymous):

whats the difference?

OpenStudy (anonymous):

sorry i was gone for over an hour

hero (hero):

2(x+y) = add x and y, then multiply by two 2x + y = multiply 2 and x, then add y

hero (hero):

Hint: Recall Order of Operations